### v1

# Distilling Diffusion Models to Efficient 3D LiDAR Scene Completion

Abstract
Diffusion models have been applied to 3D LiDAR scene completion due to their strong training stability and high completion quality. However, the slow sampling speed limits the practical application of diffusion-based scene completion models since autonomous vehicles require an efficient perception of surrounding environments. This paper proposes a novel distillation method tailored for 3D LiDAR scene completion models, dubbed ScoreLiDAR, which achieves efficient yet high-quality scene completion. ScoreLiDAR enables the distilled model to sample in significantly fewer steps after distillation. To improve completion quality, we also introduce a novel Structural Loss, which encourages the distilled model to capture the geometric structure of the 3D LiDAR scene. The loss contains a scene-wise term constraining the holistic structure and a point-wise term constraining the key landmark points and their relative configuration. Extensive experiments demonstrate that ScoreLiDAR significantly accelerates the completion time from 30.55 to 5.37 seconds per frame ($>$5$\times$) on SemanticKITTI and achieves superior performance compared to state-of-the-art 3D LiDAR scene completion models. Our code is publicly available on https://github.com/happyw1nd/ScoreLiDAR.
Figure 1: A demonstration of the LiDAR scene completion examples. Given a sparse LiDAR scan in (a), the model aims to recover the ground-truth dense scene as in (b). In these examples, scans are from SemanticKITTI [1] and KITTI360 [17] dataset. In both cases, LiDiff [24], a SOTA LiDAR scene completion method, requires about 30 seconds as in (c). In comparison, our proposed ScoreLiDAR takes only about 5 seconds in (d), achieving over 5x speedup with improved completion quality indicated by lower Chamfer Distance (CD).
1 Introduction
Recognizing the surrounding environment accurately and efficiently using onboard sensors is crucial for the safe operation of autonomous vehicles [16, 15]. Among different types of sensors, 3D LiDAR has become one of the most widely adopted sensors due to its broader detection range and higher detection precision [24, 21]. However, driving scenarios are often complex, and the 3D point clouds collected by LiDAR are typically sparse, particularly in occluded areas [34, 14]. This sparsity causes a decline in the ability to understand 3D scenes [24, 3]. Thus, inferring and completing sparse 3D LiDAR scenes is necessary to provide a dense and more comprehensive scene representation.
Figure 2: A visualization of LiDAR scene completion performances with different models on SemanticKITTI [1] dataset. Our proposed ScoreLiDAR with 8 sampling steps performs better than LiDiff [24] with 50 steps, as shown by a lower Chamfer Distance yet with less time cost. Generally, ScoreLiDAR achieves better scene completion performance and speed trade-off.
Due to the advantages of strong training stability and high-generation quality, existing works utilize diffusion models to complete the 3D LiDAR scenes and achieve outstanding results [24, 23]. However, the diffusion model often requires multiple network iterations to obtain a dense, complete, and high-quality LiDAR scene, which is time-consuming [50, 12]. Autonomous vehicles require fast and efficient perception and recognition of surrounding environments, so the slow sampling speed limits the practical application of diffusion models. Although existing works have proposed acceleration methods for diffusion models [33, 29, 19, 46, 36], 3D LiDAR scenes contain complex geometric structure information, directly applying existing acceleration methods may lead to degraded local details and reduced realism in the completed scene.
In this work, we propose ScoreLiDAR, a novel distillation method tailored for 3D LiDAR scene completion diffusion models, which enables efficient and high-quality scene completion. Variational Score Distillation (VSD) [37] uses a pre-trained diffusion model to calculate a distribution matching loss for training a student model, which has achieved impressive results. Inspired by this, ScoreLiDAR adapts and expands VSD for effective distillation of the pre-trained 3D LiDAR scene completion diffusion model. Moreover, we introduce a Structural Loss to ensure the stability of training and improve the final performance. The structural loss contains a scene-wise term constraining the holistic structure and a point-wise term constraining the key landmark points and their relative configuration, which helps the student model to capture the geometric structure information of 3D LiDAR scenes and achieve high-quality completion. We compared the proposed ScoreLiDAR with the state-of-the-art (SOTA) LiDAR scene completion models. Extensive experiments demonstrate that ScoreLiDAR can effectively accelerate the sampling speed of LiDAR scene completion diffusion models while achieving optimal scene completion quality, as shown in Fig. 1 and Fig. 2.
Our contribution can be summarized as follows: (1) We propose ScoreLiDAR, a novel distillation method tailored for diffusion-based 3D LiDAR scene completion models, which achieves efficient scene completion. (2) We introduce a Structural Loss to effectively capture the geometric structure information of 3D point clouds during the distillation process, which ensures high-quality scene completion. (3) Extensive experiments show that ScoreLiDAR enables fast and efficient scene completion while achieving optimal generation quality compared to the existing models.
2 Related work
3D LiDAR Scene completion
3D LiDAR scene completion refers to recovering a complete scene from sparse, incomplete LiDAR scan in applications such as autonomous driving [41, 34]. Current mainstream LiDAR scene completion methods include depth completion-based and Signed Distance Field (SDF)-based approaches. Depth completion-based methods aim to recover dense depth maps from sparse depth measurements [42, 8, 38]. These methods typically leverage deep learning techniques [4, 7] and can also incorporate guidance from RGB images to achieve higher-quality completion results [49, 28, 47]. SDF-based methods represent scenes as voxel grids, with the core idea of using signed distance fields to complete sparse LiDAR scene [34, 14]. These methods are constrained by voxel resolution, making them prone to losing details within the scene [24, 6]. In addition, some methods introduce semantic information to enhance LiDAR scene completion [27, 39]. These methods can generate dense and complete scenes while providing semantic labels for each point, leading to broader application potential [43, 35].
Diffusino-based 3D LiDAR scene completion
Due to the strong training stability and high generation quality of diffusion models, many methods leverage diffusion models for LiDAR scene completion tasks [24, 3, 23, 25, 13]. The work of Lee et al[13] is the first to apply diffusion models at the scene scale for LiDAR scene completion, enabling the generation of realistic scenes conditioned on partial observations from sparse point clouds. Similarly, R2DM [23] utilizes diffusion models based on distance and reflectance intensity image representations to generate various high-fidelity 3D LiDAR scenes. LiDiff [24] indicates that adding noise to point cloud data at the scene scale leads to a loss of detail. Therefore, LiDiff proposes operating directly on individual points and redefines the noise schedule and denoising processes to generate scenes with richer detail. Based on LiDiff, DiffSSC [3] further performs semantic scene completion tasks by implementing denoising and segmentation separately in both the point and semantic spaces. Moreover, LiDMs [25] constructs the pipeline from the perspectives of modal realism, geometric realism, and object realism, achieving generation under different conditions.
However, due to the inherently slow sampling process of diffusion models, the inference of these diffusion-based 3D LiDAR scene completion models is relatively slow. This limitation makes it challenging to achieve fast and efficient perception as required in autonomous vehicle applications.
3 Preliminary
3.1 Brief introduction of diffusion models
The diffusion models have two processes: forward diffusion and reverse denoising process [9, 30]. In the forward diffusion process, given the data $\boldsymbol{x}^{0}\sim q(\boldsymbol{x})$ from the training distribution, the diffusion model adds different scales of noise to $\boldsymbol{x}^{0}$ according to different timesteps $t\in[1,T]$ to obtain noisy data $\{\boldsymbol{x}^{1},\boldsymbol{x}^{2},\ldots,\boldsymbol{x}^{T}\}$. When $T$ is large enough, $\boldsymbol{x}^{T}$ approaches to standard Gaussian distribution, namely, $q(x^{T})\approx\mathcal{N}(0,I)$. This process is parameterized by a series of predefined noise factors $\beta^{t}$. By defining $\alpha^{t}=1-\beta^{t}$, the diffusion process is expressed as [9]:
$\boldsymbol{x}^{t}=\sqrt{\bar{\alpha}^{t}}\boldsymbol{x}^{0}+\sqrt{1-\bar{\alpha}^{t}}\boldsymbol{\epsilon}^{t}$ (1)
Here $\bar{\alpha}^{t}=\prod_{t=1}^{T}\alpha^{t}$, $p(\boldsymbol{x}^{t}\mid\boldsymbol{x}^{0})=\mathcal{N}(\sqrt{\bar{\alpha}^{t}},(1-\bar{\alpha}^{t})I)$.
During the training, the diffusion model tries to predict the added noise at different timesteps $t$. Given the input $\boldsymbol{x}^{0}$ and the condition $c$ (optional), the noisy data $\boldsymbol{x}^{t}$ can be calculated by Eq. 1. The diffusion model $\boldsymbol{\epsilon}_{\theta}$ predicts the noise according to $\boldsymbol{x}^{t},c,t$ and is then optimized by calculating the $\ell_{2}$ loss between the predicted and the real noise.
$\mathcal{L}_{DM}=\mathbb{E}_{t,\epsilon}\left[\|\boldsymbol{\epsilon}^{t}-\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}^{t},c,t)\|^{2}\right]$ (2)
Here $\theta$ is the trainable parameter of $\boldsymbol{\epsilon}_{\theta}$.
In the reverse denoising process, the diffusion model starts from the timestep $T$ and progressively removes the predicted noise until a generated sample is obtained. The process of denoising $\boldsymbol{x}^{t}$ to obtain $\boldsymbol{x}^{t-1}$ can be written as:
$\boldsymbol{x}^{t-1}=\frac{1}{\sqrt{\alpha^{t}}}\left(\boldsymbol{x}^{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}^{t}}}\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}^{t},c,t\right)\right)+\sigma^{t}\boldsymbol{z}$ (3)
Here $z\sim\mathcal{N}(0,I)$. In this process, the number of required inference steps varies depending on different sampling methods. For instance, DDPM [9] requires 1000 timesteps, while DDIM [30] and DPM solver [18] can reduce this to 100 timesteps and 20 timesteps, respectively.
3.2 3D LiDAR scene completion diffusion models
The 3D LiDAR scene completion diffusion models take the incomplete scan $\mathcal{P}=\{\boldsymbol{p}_{1},\boldsymbol{p}_{2},...,\boldsymbol{p}_{N}\}$ and try to generate the complete scene $\mathcal{G}^{0}=\{\boldsymbol{p}^{0}_{1},\boldsymbol{p}^{0}_{2},...,\boldsymbol{p}^{0}_{M}\}$. Given the input LiDAR scan $\mathcal{P}$ and ground truth $\mathcal{G}$, a diffusion model can be trained to perform 3D LiDAR scene completion. The noisy scene $\mathcal{G}^{t}$ at timestep $t$ is calculated from the ground truth $\mathcal{G}$ at point level [24, 3],
$\boldsymbol{p}_{m}^{t}=\sqrt{\bar{\alpha}^{t}}\boldsymbol{p}_{m}+\sqrt{1-\bar{\alpha}^{t}}\epsilon^{t},\forall\boldsymbol{p}_{m}\in\mathcal{G}$ (4)
Here $\mathcal{G}^{t}=\{\boldsymbol{p}^{t}_{1},\boldsymbol{p}^{t}_{2},...,\boldsymbol{p}^{t}_{M}\}$. Because the LiDAR point cloud is sparse, the noisy data retains very little information about the original data. To generate more realistic point cloud scenes, the LiDAR scan $\mathcal{P}$ can be used as a condition of the diffusion model [24]. In this case, the training loss of the diffusion model is given by:
$\mathcal{L}_{DM}=\mathbb{E}_{t,\epsilon}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\mathcal{G}^{t},\mathcal{P},t\right)\right\|^{2}\right]$ (5)
Then, as described in Sec. 3.1, the completed scene ${\mathcal{G}}^{0}$ can be generated by progressive denoising from $\mathcal{G}^{T}$. Because the scale of the LiDAR scene is large and the data range is different across different point cloud axes, directly normalizing the entire dataset compresses the data into a smaller range, which potentially leads to the loss of critical details [24, 3]. To solve this issue, LiDiff [24] modifies the diffusion process by adding a local noise offset to each point $\boldsymbol{p}_{m}$, gradually perturbing the point cloud at each timestep. For Eq. 1, $\boldsymbol{x}_{0}$ is set to 0, and $\boldsymbol{x}_{t}$ is added to each point $\boldsymbol{p}_{m}$,
$\boldsymbol{p}_{m}^{t}=\boldsymbol{p}_{m}+\left(\sqrt{\bar{\alpha}_{t}}\mathbf{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon^{t}}\right)=\boldsymbol{p}_{m}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon^{t}}$ (6)
Due to this special case, $\mathcal{G}^{T}$ cannot directly start from standard Gaussian noise in the sampling process. Instead, the LiDAR scan $\mathcal{P}$ is used to obtain $\mathcal{G}^{T}$ [24]. Firstly, given the initial incomplete scan $\mathcal{P}$, the number of the point clouds is increased by duplicating the original points $K$ times and getting the pseudo dense scan $\mathcal{P}^{*}=\{\boldsymbol{p}^{*}_{1},\boldsymbol{p}^{*}_{2},\ldots,\boldsymbol{p}^{*}_{M}\}$, where we assume $M=KN$. Then, we calculate the noisy point cloud $\mathcal{P}^{T}$ by Eq. 6. As $\mathcal{P}^{T}$ is noisy enough, it can be regarded as $\mathcal{G}^{T}$ during the training. After that, a step-by-step denoising process is applied to obtain the completed scene $\mathcal{G}^{0}$.
Figure 3: The overall structure of ScoreLiDAR. (1) The student model generates the completed scene based on the sparse scan. (2) The sparse scan and noisy completed scene are input to $\boldsymbol{\epsilon}_{\theta}$ and $\boldsymbol{\epsilon}_{\phi}$. (3) The predicted score of $\boldsymbol{\epsilon}_{\theta}$ and $\boldsymbol{\epsilon}_{\phi}$ are used to calculated the KL divergence. (4) Structural loss is calculated based on the completed scene and the ground truth. (5) The total loss is calculated with KL divergence and structural loss. (6) The student model is optimized according to the total loss. (7) The diffusion model $\boldsymbol{\epsilon}_{\phi}$ is updated with the completed scene.
4 Method
Our goal is to distill a pre-trained 3D LiDAR scene completion diffusion model into a student model with significantly fewer sampling steps, enabling efficient and high-quality scene completion. Firstly, we introduce the distillation method tailored for 3D LiDAR scene completion diffusion models in Sec. 4.1. Then, we introduce the structural loss to improve the distillation process with both scene-wise loss and point-wise loss in Sec. 4.2. Finally, we describe the optimization procedure of ScoreLiDAR in Sec. 4.3.
4.1 Distillation for 3D LiDAR scene completion
[…]
$\min_{\eta}D_{KL}\left(p^{0}_{G}\left(\mathcal{G}^{0};\eta\right)\|q^{0}\left(\mathcal{G}^{0}\right)\right)$ (7)
[…]
$\min_{\eta}\mathcal{L}_{KL}=\mathbb{E}_{t,\epsilon}D_{KL}\left(p^{t}_{G}\left(\mathcal{G}^{t}\right)\|q^{t}\left(\mathcal{G}^{t}\right)\right)$ (8)
[…]
$\nabla_{\eta}D_{\mathrm{KL}}\left(p_{G}^{t}\left(\mathcal{G}^{t}\right)\|q^{t}\left(\mathcal{G}^{t}\right)\right)$ (9)
$=\mathbb{E}_{t,\epsilon}\left[\nabla_{\mathcal{G}^{t}}\log p_{G}^{t}\left(\mathcal{G}^{t}\right)-\nabla_{\mathcal{G}^{t}}\log q^{t}\left(\mathcal{G}^{t}\right)\right]\frac{\partial\mathcal{G}^{t}}{\partial\eta}$
We use the pre-trained diffusion model $\boldsymbol{\epsilon}_{\theta}$ to approximate $\nabla_{\mathcal{G}^{t}}\log q^{t}\left(\mathcal{G}^{t}\right)$ with $\nabla_{\mathcal{G}^{t}}\log q^{t}\left(\mathcal{G}^{t}\right)\approx-\frac{\boldsymbol{\hat{\epsilon}}}{\sqrt{1-\bar{\alpha}^{t}}}$ [32]. Similarly, the score $\nabla_{\mathcal{G}^{t}}\log p_{G}^{t}\left(\mathcal{G}^{t}\right)$ can be approximated by another auxiliary diffusion model $\boldsymbol{\epsilon}_{\phi}$. Then, with the simplification as in [9], the $\mathcal{L}_{KL}$ is estimated by
$\mathcal{L}_{KL}\approx\mathbb{E}_{t,\epsilon}\left[\|\boldsymbol{\epsilon}_{\theta}\left(\mathcal{G}^{t},\mathcal{P},t\right)-\boldsymbol{\epsilon}_{\phi}\left(\mathcal{G}^{t},\mathcal{P},t\right)\|^{2}_{2}\right]$ (10)
Thus, the gradient in Eq. 9 is approximated by
$\nabla_{\eta}D_{\mathrm{KL}}\left(p_{G}^{t}\left(\mathcal{G}^{t}\right)\|q^{t}\left(\mathcal{G}^{t}\right)\right)$ (11)
$\approx\mathbb{E}_{t,\epsilon}\left[\boldsymbol{\epsilon}_{\theta}\left(\mathcal{G}^{t},\mathcal{P},t\right)-\boldsymbol{\epsilon}_{\phi}\left(\mathcal{G}^{t},\mathcal{P},t\right)\right]\frac{\partial\mathcal{G}^{t}}{\partial\eta}$
[…]
4.2 Structural loss
Although the distillation process in Sec. 4.1 is highly effective in training models [19, 50], we found that directly applying it to LiDAR scene completion diffusion models leads to loss of local details and reduced realism in the completed scene. This is because the point cloud in LiDAR scenes often includes complex geometric information that is not explicitly captured by diffusion models. Thus, we introduce a structural loss to further refine the distillation process and improve the completion quality. This structural loss includes scene-wise loss and point-wise loss and can help the student model effectively capture geometric structure information of the 3D point clouds.
Scene-wise loss.
In the distillation process mentioned in Sec. 4.1, the gradient $\nabla_{\eta}D_{\mathrm{KL}}$ in Eq. 11 is well-defined when $t\gg 0$, i.e. the generated samples are totally disturbed by Gaussian noise. However, $\nabla_{\eta}D_{\mathrm{KL}}$ becomes unreliable when $t$ is small [50, 48]. This is because the student model often generates subpar results at the early stage due to the complexity of the point cloud data. It is easy for the noisy generated samples to lie outside the training distribution of the teacher model, causing the unreliable network prediction of the teacher model [50, 48].
To solve this issue, we introduce the scene-wise loss, which minimizes the distance between the ground truth scene $\mathcal{G}$ and the completed scene $\mathcal{G}^{0}$. Concretely, we calculate the scene-wise loss by
$\mathcal{L}_{scene}=\mathbb{E}_{t,\epsilon}\left[\|\mathcal{G}-\mathcal{G}^{0}\|^{2}_{2}\right]$ (12)
This loss calculates the mean squared error between each point in the generated scene and its closest corresponding point in the ground truth. It helps the student model capture the holistic structure, which prevents the optimization direction from deviating in the early stages and enhances training stability. The scene-wise loss also enables the generated scenes to be closer to the ground truth globally, thereby enhancing the completion quality and fidelity.
Point-wise loss.
As seen in Eq. 11, the distillation process only constrains the overall distribution of the completed scene, ignoring the relative positions between different points. Directly using the gradient in Eq. 11 to optimize the student model may lead to loss of local details.
Thus, we introduce the point-wise loss to capture the relative structural information between different points in the 3D LiDAR scene. The point-wise loss calculates the difference between the inter-point distance matrices of the completed scene and the ground truth. Due to the large number of points in the scene, calculating the distance matrix for all points is computationally intensive. Therefore, we select $n$ key points to compute the distance matrix with $n\ll M$. Based on the local geometric features of each point, we choose key points that are critical for representing the structure of the 3D LiDAR scene. For each point $\boldsymbol{p}^{0}_{i}$ in the completed scene $\mathcal{G}^{0}$, we find its K-nearest neighbor, denoted as the set $\mathcal{K}_{i}$. Then we select the key points by calculating their curvature $\kappa_{i}$. The specific steps are as follows:
  • Calculate the centroid $\bar{\boldsymbol{p}}^{0}_{i}$ of the neighborhood $\mathcal{K}_{i}$
    $\bar{\boldsymbol{p}}^{0}_{i}=\frac{1}{K}\sum_{\boldsymbol{p}^{0}_{j}\in\mathcal{K}_{i}}\boldsymbol{p}^{0}_{j}$ (13)
  • Calculate the neighborhood covariance matrix $\mathcal{C}_{i}$ for $\bar{\boldsymbol{p}}^{0}_{i}$
    $\mathcal{C}_{i}=\frac{1}{K}\sum_{\boldsymbol{p}^{0}_{j}\in\mathcal{K}_{i}}(\boldsymbol{p}^{0}_{j}-\bar{\boldsymbol{p}}^{0}_{i})(\boldsymbol{p}^{0}_{j}-\bar{\boldsymbol{p}}^{0}_{i})^{T}$ (14)
  • Perform eigen-decomposition on the covariance matrix $\mathcal{C}_{i}$ to obtain the eigenvalues $\lambda_{1}<\lambda_{2}<\cdots<\lambda_{m}$.
  • Curvature $\kappa_{i}$ can be calculated using the eigenvalues
    $\kappa_{i}=\frac{\lambda_{1}}{\sum_{j=1}^{m}\lambda_{j}}$ (15)
A larger curvature $\kappa_{i}$ indicates greater local shape variation. Those points with great local variation are typically located at corners, edges, or endpoints, which tend to shape the main structure of the scene. Therefore, the top $n$ points with the highest curvature values are selected as key points.
Given the ground truth $\mathcal{G}=\{\boldsymbol{p}_{1},\boldsymbol{p}_{2},...,\boldsymbol{p}_{M}\}$, we select $n$ key points ($n\ll M$) from the point cloud to construct the $n\times n$ distance matrix $\mathcal{D}$. The $d_{ij}\in\mathcal{D}$ represents the Euclidean distance between the point $i$ and $j$. Then, for completed scene $\mathcal{G}^{0}$, we select $n$ points that are closest to the key points in $\mathcal{G}$ as the corresponding key points to obtain the distance matrix $\mathcal{D}_{G}$. Thus, the point-wise loss is calculated by
$\mathcal{L}_{point}=\mathbb{E}_{t,\epsilon}\left[\|\mathcal{D}-\mathcal{D}_{G}\|^{2}_{2}\right]$ (16)
The point-wise loss can help the student model capture the relative configuration of key points and further enhance the geometric accuracy and detail retention of the completed scene. This ensures that key objects like cars, traffic cones, and walls are better completed, which is crucial for autonomous vehicles to recognize surroundings accurately.
Structural loss.
[…]
$\mathcal{L}_{structural}=\lambda_{scene}\mathcal{L}_{scene}+\lambda_{point}\mathcal{L}_{point}$ (17)
Here $\lambda_{scene}$ and $\lambda_{point}$ are the weight of scene-wise loss and point-wise loss.
4.3 Optimization procedure
[…]
$\mathcal{L}_{stu}=\mathcal{L}_{KL}+\mathcal{L}_{structural}$ (18)
We set $\lambda_{scene}=0.5$ and $\lambda_{point}=0.01$ unless otherwise specified. The implementation details are provided in Sec. A.3.
Model CD $\downarrow$ JSD $\downarrow$ Times (s) $\downarrow$
LMSCNet [26] 0.641 0.431 0.40
LODE [14] 1.029 0.451 1.76
MID [34] 0.503 0.470 6.42
PVD [51] 1.256 0.498 -
LiDiff [24] 0.434 0.444 30.38
LiDiff (Refined) [24] 0.375 0.416 30.55
ScoreLiDAR 0.406 0.425 5.16
ScoreLiDAR (Refined) 0.342 0.399 5.37
[…]
Model CD $\downarrow$ JSD $\downarrow$ Times (s) $\downarrow$
LMSCNet [26] 0.979 0.496 -
LODE [14] 1.565 0.483 -
MID [34] 0.637 0.476 -
LiDiff [24] 0.564 0.459 29.18
LiDiff (Refined) [24] 0.517 0.446 29.43
ScoreLiDAR 0.472 0.444 4.98
ScoreLiDAR (Refined) 0.452 0.437 5.14
Table 2: The completion performance on the KITTI-360 dataset. The meaning of notations is the same as those in Tab. 1.
Model SemanticKITTI KITTI360
CD $\downarrow$ JSD $\downarrow$ CD $\downarrow$ JSD $\downarrow$
ScoreLiDAR (Refined) 0.342 0.399 0.452 0.437
w/o Structural Loss 0.419 0.430 0.549 0.445
Table 3: Ablation study of the structural loss.
Model CD $\downarrow$ JSD $\downarrow$ Time (s) $\downarrow$
LiDiff (50 steps) [24] 0.434 0.444 30.38
LiDiff (50 steps Refined) [24] 0.375 0.416 30.55
LiDiff (8 steps) [24] 0.447 0.432 5.69
LiDiff (8 steps Refined) [24] 0.411 0.406 5.92
ScoreLiDAR (8 Steps Refined) 0.342 0.399 5.37
ScoreLiDAR (4 Steps Refined) 0.326 0.386 3.23
ScoreLiDAR (2 Steps Refined) 0.403 0.379 1.85
ScoreLiDAR (1 Steps Refined) 0.548 0.384 1.10
Table 4: Ablation study of different sampling steps on the SemanticKITTI dataset.
5 Experiment
In this part, we conduct a series of experiments to evaluate the effectiveness of the proposed ScoreLiDAR. We compare ScoreLiDAR with advanced models including LMSCNet [26], LODE [14], MID [34], PVD [51] and LiDiff [24]. We first evaluate the performance of ScoreLiDAR in scene completion tasks (Sec. 5.1). Secondly, we present the results of ablation studies showing the effectiveness of the structural loss and the performances of ScoreLiDAR given different sampling steps (Sec. 5.2). Finally, we further evaluate ScoreLiDAR with the qualitative analysis ( Sec. 5.3).
Figure 4: Qualitative results on KITTI-360. ScoreLiDAR achieves better completion than LiDiff [25] with fewer sampling steps.
5.1 Scene completion
We validate ScoreLiDAR on SemanticKITTI [1] and KITTI-360 [17] datasets. The existing SOTA LiDAR scene completion model LiDiff [24] is chosen as the teacher model. The student model shares the network architecture with the teacher model and is initialized by the teacher model. Moreover, we also use the refinement network in LiDiff [24] to refine the completed scene generated by the student model. We calculate the Chamfer Distance (CD) and the Jensen-Shannon Divergence (JSD) to evaluate the similarity between the completed scene and the ground truth. The smaller the value of CD and JSD, the closer the completed scene is to the ground truth.
Tab. 1 shows the quantitative results on SemanticKITTI. ScoreLiDAR achieves the optimal performance compared to the existing models on both metrics. Compared to the SOTA method LiDiff [24] with refinement, which takes 30.55 seconds to complete a scene, our proposed ScoreLiDAR completes a scene in just 5.47 seconds (fivefold speedup) yet with $8\%$ improvement in CD and $4\%$ in JSD. Although LMSCNet [26] and LODE [14] have faster completion speeds, their completion quality is significantly lower. Notably, the performance of ScoreLiDAR outperforms the teacher model LiDiff [24]. This is because ScoreLiDAR introduces a structural loss with scene-wise term and point-wise term, enabling the student model to effectively capture geometric structure information within LiDAR point cloud data during training. Results on KITTI-360 are shown in Tab. 2. ScoreLiDAR also achieves optimal performance compared to existing models. ScoreLiDAR also boasts a fivefold speedup with $12\%$ improvement in CD and $2\%$ in JSD compared to LiDiff [24].
5.2 Ablation study
In this part, we conduct the ablation study to verify the effectiveness of the structural loss in the training of the proposed ScoreLiDAR. We compared the scene completion performances of the proposed ScoreLiDAR with a variant that does not incorporate structural loss. The results are shown in Tab. 3. The results show that the variant without structural loss exhibits lower performance in scene completion on both datasets. However, after considering the structural loss, the performance of ScoreLiDAR improves significantly, which achieves better performance on both metrics. This supports our discussion in Sec. 4.3, incorporating structural loss enables the student model to capture the geometric structure feature of 3D point clouds, thereby facilitating the effective distillation of the student model.
Furthermore, we compared the scene completion performance of ScoreLiDAR with different sampling steps, and the results are shown in Tab. 4. It can be observed that as the sampling steps decrease from 8 to 1, the time required for ScoreLiDAR to complete a scene also decreases, with single-step sampling allowing a scene to be completed in only 1.1 seconds. With 8-step and 4-step sampling, ScoreLiDAR performs better on both metrics than LiDiff. Both metrics decay at 2-step and 1-step sampling, but in JSD ours still performs better than LiDiff. In summary, although the quality of scene completion decreases as the sampling steps are reduced, it still maintains performance comparable to or better than the existing model, achieving better performance and speed trade-off as in Fig. 2.
Figure 5: The qualitative analysis of structural loss. The bar chart shows the distribution of distances between corresponding points in the completed and ground truth scenes. A higher number of points with smaller distances demonstrates that the completed scene is closer to the ground truth. The heatmap represents the difference in distance matrices between the completed scene and the ground truth scene. Smaller values on the heatmap indicate that the completed scene is closer to the ground truth.
5.3 Qualitative analysis
Fig. 4 shows the completed scenes by our proposed ScoreLiDAR and LiDiff [24] on KITTI-360. ScoreLiDAR achieves completion results with higher quality and greater fidelity. We can see that LiDiff [24] nearly fails at 8 steps. Although LiDiff [24] achieves decent completion at 50 steps, there are some missing areas on the lower and right sides. In contrast, ScoreLiDAR reaches optimal completion including clearer objects such as cars, traffic cones, and other scene elements with only 8 steps, which is closer to the ground truth, especially after refinement.
To further demonstrate the effectiveness of ScoreLiDAR and the structural loss, we calculate the distance between the points in the completed scene and their corresponding points in the ground truth to evaluate the overall difference. We display the calculated results in bar chat in Fig. 5. ScoreLiDAR has the highest number of points with smaller distances to their corresponding points in the ground truth. The results show that the scenes completed by ScoreLiDAR are closer to the ground truth overall, demonstrating higher fidelity. Moreover, we selected 36 corresponding key points from the ground truth and the completed scene using the method described in Sec. 4.2 and calculated the point distance matrices $\mathcal{D}$ and $\mathcal{D_{G}}$. We then visualized the difference between $\mathcal{D}$ and $\mathcal{D_{G}}$ as a heatmap. As shown in Fig. 5, on both datasets, the difference of point distance matrix between the completed scene of LiDiff [24] and the ground truth is the largest, followed by the ScoreLiDAR variant without the structural loss and the smallest difference is achieved by ScoreLiDAR. This also indicates that the scene completed by ScoreLiDAR is closer to the ground truth.
We also conduct a user study to evaluate the performance of ScoreLiDAR. We used ScoreLiDAR and LiDiff [24] to complete scenes based on the same input scans and asked users to choose the scene they believed was closer to the ground truth. ScoreLiDAR received a $65\%$ user preference over LiDiff [24]. This indicates that the detail and fidelity of the scenes completed by ScoreLiDAR more closely resemble the ground truth for most users. The details of the user study are shown in Sec. D.4.
6 Conclusion
Summary.
This paper proposes ScoreLiDAR, a novel distillation method tailored for 3D LiDAR scene completion diffusion models. The distilled model enables efficient LiDAR scene completion. By introducing the structural loss with scene-wise term and point-wise term, ScoreLiDAR trains the student model to effectively capture the holistic structure and the relative configuration of key points and achieve efficient and high-quality scene completion.
Limitations.
While ScoreLiDAR achieves efficient, high-quality LiDAR scene completion, its performance is constrained by the teacher model. As the performance of the teacher model improves, so does the capability of the student model. Moreover, the ability to perform semantic scene completion is also determined by the teacher model. If the teacher model can complete the semantic scene, the distilled student model will also be able to follow the teacher. Thus, further exploration is required to find a more effective method to improve the training process [45, 44] of ScoreLiDAR and avoid the limitations of the teacher model, achieving more efficient semantic LiDAR scene completion.
Appendix A Experiment protocol
A.1 Dataset setup
SemanticKITTI [1] dataset is a large-scale benchmark for 3D semantic segmentation in autonomous driving, extending the KITTI Odometry dataset with dense semantic annotations for over 43,000 LiDAR scans. It provides labels for 25 classes, such as “car,” “road,” and “building,” capturing diverse urban and rural scenes. The SemanticKITTI dataset consists of 22 sequences, where sequences 00-10 are densely annotated for each scan, enabling tasks such as semantic segmentation and semantic scene completion using sequential scans. Sequences 11-21 serve as the test set, showcasing diverse and challenging traffic situations and environment types to evaluate model performance in real-world autonomous driving scenarios. SemanticKITTI is widely used in research and serves as a critical resource for advancing LiDAR-based perception systems.
KITTI-360 [17] dataset is a comprehensive benchmark for 3D scene understanding in autonomous driving, capturing 360-degree panoramic imagery and 3D point clouds across diverse urban environments. It includes over 73 km of driving data with dense semantic annotations for both 2D (images) and 3D (point clouds), covering categories like “vehicles,” “buildings,” and “vegetation.” KITTI-360 provides high-resolution sensor data, including LiDAR, GPS/IMU, and stereo camera recordings, making it ideal for tasks such as 3D semantic segmentation, panoptic segmentation, and mapping in real-world driving scenarios.
Figure 6: Comparison of scene completion details between ScoreLiDAR and the SOTA model LiDiff [24]. The magnified images are enlarged views of the regions corresponding to the boxes in the completed scene images. Compared to LiDiff, ScoreLiDAR better completes the details of vehicles, making it closer to the ground truth scene.
A.2 Evaluation metrics
Chamfer Distance (CD)
[2] is a metric used to measure the similarity between two sets of points, often employed for evaluating the quality of generated point clouds or geometric shapes. For two point sets $P$ and $Q$, the Chamfer Distance is defined as:
$CD(P,Q)=\frac{1}{|P|}\sum_{p\in P}\min_{q\in Q}\|p-q\|^{2}+\frac{1}{|Q|}\sum_{q\in Q}\min_{p\in P}\|q-p\|^{2}$ (19)
The first term calculates the average squared distance from each point in $P$ to its nearest neighbour in $Q$. The second term calculates the average squared distance from each point in $Q$ to its nearest neighbour in $P$. Chamfer Distance evaluates how well two point sets approximate each other by considering their nearest neighbour distances in both directions. CD effectively captures local geometric features and exhibits strong robustness in local shape matching, which is commonly used in evaluating the matching and reconstruction of 3D point clouds.
Jensen-Shannon Divergence (JSD)
[20] is a symmetric measure of similarity between two probability distributions. It is a variation of the Kullback-Leibler (KL) divergence and is widely used in information theory, statistics, and machine learning. Given two probability distributions $P$ and $Q$ over the same domain, JSD is defined as:
$JSD(P\|Q)=\frac{1}{2}KL(P\|M)+\frac{1}{2}KL(Q\|M)$ (20)
Here $M=\frac{1}{2}(P+Q)$ is the average distribution, and $KL(P||M)$ is the Kullback-Leibler divergence.
[…]
A.3 Implementation details
We choose the pre-trained LiDiff [24] model as the teacher model $\boldsymbol{\epsilon_{\theta}}$, the student model $G_{stu}$ and the auxiliary diffusion model $\boldsymbol{\epsilon}_{\phi}$ shares the same network architecture as the teacher model and are initialized by the teacher model. The ScoreLiDAR is trained on SemanticKITTI dataset. The pre-trained diffusion model is provided by the official release of LiDiff [24].
For optimization, we use the Stochastic Gradient Descent (SGD) optimizer with the default parameters. The learning rate is set to $53e-$ 5 and the batch size is set to $1$. The training ratio between the student model and the auxiliary diffusion model is maintained at $1:1$. To reduce computational costs, when calculating the point-wise loss, we first randomly select $\frac{1}{10}$ of the points from the ground truth scene. Then, following the proposed method, we select the top $\frac{1}{3}$ points with the highest curvature from these points as the key points to calculate the distance matrix. That is, the final number of key points is $\frac{1}{30}$ of the total number of points in the ground truth scene. When calculating the $K$-nearest neighbours, we set $K=180$. The weights of scene-wise loss $\lambda_{scene}$ and the point-wise loss $\lambda_{point}$ are set to $0.5$ and $0.01$, respectively. ScoreLiDAR requires only 50 iterations to achieve convergence, taking approximately 10 minutes on a single A40 GPU, which is highly efficient. Our code will be publicly available soon.
Appendix B Discussion
Firstly, we discuss the significance of this study. For autonomous vehicles, accurately recognizing and perceiving their surrounding environment during operation is critical [16, 15]. This is particularly important for identifying objects that may affect the vehicle’s movement, such as other vehicles, pedestrians, traffic cones, and signposts [11]. Accurate and efficient recognition of these objects is essential for the safe operation of autonomous vehicles. However, the scan data obtained by onboard LiDAR as shown in Fig. 6 (a) is sparse [34, 14], and it is difficult to identify key objects such as vehicles from the magnified regions of the sparse scan. However, from the ground truth scene in Fig. 6 (b), it is evident that these regions contain moving vehicles. Autonomous vehicles cannot obtain sufficient information about the driving environment from these sparse LiDAR scans [25, 13]. Therefore, it is necessary to use appropriate methods to complete the sparse LiDAR scans.
LiDiff [24] uses DDPM [9] models to complete 3D LiDAR scenes, achieving impressive results. However, due to the inherent characteristics of diffusion models, LiDiff [24] requires approximately 30 seconds to complete a single scene, limiting its applicability in autonomous vehicles. In contrast, the proposed ScoreLiDAR can complete a scene in almost 5 seconds, more than 5 times faster than LiDiff [24], while achieving higher completion quality. As shown in Fig. 6 (c), in the magnified region, although LiDiff [24] completes some vehicle shapes, they are not fully reconstructed. For example, in the left region, there should be two vehicles, but the scene completed by LiDiff only contains one. In contrast, the scene completed by ScoreLiDAR features clearer and more complete vehicle structures (Fig. 6 (d)), making it closer to the ground truth. Thus, with the scenes completed by the proposed ScoreLiDAR, autonomous vehicles can more easily recognize critical objects in their driving environment, enabling safer and more effective navigation.
Moreover, ScoreLiDAR allows for different sampling steps during distillation to achieve varying completion speeds. However, due to the limitations of the teacher model, ScoreLiDAR’s completion quality decreases at smaller sampling steps. With further improvements or by replacing the teacher model for better distillation, ScoreLiDAR could complete a scene within one second, achieving real-time scene completion. This is crucial for autonomous vehicles to scan and recognize their driving environment effectively. Therefore, future work will focus on exploring this aspect.
Appendix C Additional completed scenes
Fig. 8 and Fig. 9 show additional completed scenes by the proposed ScoreLiDAR and compare them with the scenes completed by LiDiff [24].
Appendix D Additional experiment results
D.1 More ablation study of structural loss
Model SemanticKITTI KITTI360
CD $\downarrow$ JSD $\downarrow$ CD $\downarrow$ JSD $\downarrow$
ScoreLiDAR 0.342 0.399 0.452 0.437
w/o Point-wise loss 0.351 0.414 0.485 0.486
w/o Scene-wise loss 0.367 0.422 0.477 0.451
w/o Structural Loss 0.419 0.430 0.549 0.445
Table 5: Ablation study of the scene-wise and point-wise loss. The metrics refer to the performance with refinement. Colors denote the 1st, 2nd, and 3rd best-performing model.
ScoreLiDAR SemanticKITTI KITTI360
CD $\downarrow$ JSD $\downarrow$ CD $\downarrow$ JSD $\downarrow$
$\lambda_{scene}=0.5,\lambda_{point}=0.01$ 0.342 0.399 0.452 0.437
$\lambda_{scene}=0.05,\lambda_{point}=0.01$ 0.354 0.409 0.494 0.457
$\lambda_{scene}=0.5,\lambda_{point}=0.1$ 0.358 0.419 0.539 0.474
Table 6: Ablation study of the different weights of the scene-wise and point-wise loss. The first row refers to the default configuration of the ScoreLiDAR. The metrics refer to the performance with refinement.
Model SemanticKITTI KITTI360
CD $\downarrow$ JSD $\downarrow$ CD $\downarrow$ JSD $\downarrow$
LiDiff (Refined) 0.375 0.416 0.517 0.446
w/ Structural loss 0.399 0.426 0.535 0.450
Table 7: Ablation study of training LiDiff [24] with structural loss.
To further validate the effectiveness of the structural loss, we evaluated the performance of variants trained with only point-wise loss or scene-wise loss and compared them with default ScoreLiDAR. As shown in Tab. 5, compared to the default ScoreLiDAR, the performance of variants trained with only scene-wise loss or point-wise loss decreased. However, compared to the variants without structural loss, the variants using only one type of loss still showed improved completion performance. These results confirm the effectiveness of the structural loss in the distillation process.
Additionally, we investigated the impact of different weights of scene-wise and point-wise loss on the completion quality. The results are shown in Tab. 6. It can be observed that reducing $\lambda_{scene}$ or increasing $\lambda_{point}$ leads to a decline in the performance of ScoreLiDAR but still achieves a comparable performance. This verifies the effectiveness of the proposed structural loss in improving the completion performance of the student model.
Finally, we trained LiDiff using structural loss to investigate whether structural loss can enhance the performance of LiDiff. The results are shown in Tab. 7. Training LiDiff [24] with structural loss does not result in a performance improvement. This may be because structural loss is not suitable for direct addition to the training loss of LiDiff [24], i.e. the traditional diffusion model training loss.
D.2 Ablation study of different key point number
ScoreLiDAR SemanticKITTI KITTI360
CD $\downarrow$ JSD $\downarrow$ CD $\downarrow$ JSD $\downarrow$
$n=1/30$ 0.342 0.399 0.452 0.437
$n=1/60$ 0.346 0.409 0.452 0.471
Table 8: Ablation study of different key points number. The first row refers to the default configuration of the ScoreLiDAR. The metrics refer to the performance with refinement.
As mentioned in Sec. A.3, the optimal number of the key point is set to the $\frac{1}{30}$ of the total number of points in the ground truth. To investigate the impact of different numbers of key points on the completion performance of ScoreLiDAR, we decreased the number of key points for model training and evaluated the completion performance. As shown in Tab. 8, when the number of key points decreases, the performance of ScoreLiDAR declines. This is because an insufficient number of key points causes the point-wise loss to fail in effectively capturing the relative positional information between key points, preventing the student model from learning the local geometric structure, and thereby reducing the completion quality.
D.3 Ablation study of different sampling steps on KITTI-360
We also conduct the ablation study of different sampling steps on the KITTI-360 dataset. The results are shown in Tab. 9. Similar to the results on the SemanticKITTI dataset, as the number of sampling steps decreases, the time required for ScoreLiDAR to complete a scene is reduced. Although the completion performance declines slightly, it remains comparable to that of existing SOTA models.
Model CD $\downarrow$ JSD $\downarrow$ Time (s) $\downarrow$
LiDiff (50 steps) [24] 0.564 0.549 29.18
LiDiff (50 steps Refined) [24] 0.517 0.446 29.43
LiDiff (8 steps) [24] 0.619 0.471 5.46
LiDiff (8 steps Refined) [24] 0.550 0.462 5.77
ScoreLiDAR (8 Steps) 0.452 0.437 5.14
ScoreLiDAR (4 Steps) 0.482 0.461 3.16
ScoreLiDAR (2 Steps) 0.525 0.457 1.69
ScoreLiDAR (1 Steps) 0.750 0.478 1.03
Table 9: Ablation study of different sampling steps on the KITTI-360 dataset. The metrics of ScoreLiDAR refer to the performance with refinement.
D.4 User study
The user study is conducted to verify the completion performance of ScoreLiDAR further. We first used ScoreLiDAR and the current SOTA method LiDiff [24] to complete the same 30 input LiDAR scans, resulting in 30 pairs of completed scenes. We then randomly recruited seven volunteers and guided each to evaluate the detail and fidelity of these 30 pairs of scene images, selecting the one they believed to be closer to the ground truth. The seven volunteers included five men and two women, aged 24–30, with five participants having research backgrounds related to autonomous driving or LiDAR scene completion and the remaining two participants having backgrounds related to artificial intelligence. They were given unlimited time for the evaluation, but the average completion time for all volunteers was 30 minutes.
Model User preference $\uparrow$
LiDiff [24] 35%
ScoreLiDAR 65%
Table 10: Results of user study. Our ScoreLiDAR outperforms the existing SOTA model.
The result of the user study is shown in Tab. 10. Compared to LiDiff, ScoreLiDAR received a $65\%$ user preference, surpassing the majority threshold. This indicates that, in the eyes of most users, the detail and fidelity of the scenes completed by ScoreLiDAR more closely resemble the ground truth. The results of the user study further demonstrate the effectiveness of ScoreLiDAR in LiDAR scene completion.
D.5 Visualization of key points
To validate the feasibility of our proposed key point selection method, we visualized the selected key points in the ground truth scene. As shown in Fig. 7, the red key points are mostly distributed on walls, traffic cones, cars, and corners, while smooth areas such as the road surface have no key points. These key points are crucial for expressing the details of 3D LiDAR scenes. Selecting these points to compute the point-wise loss allows the student model to more easily capture the relative configuration information between key points, thereby better completing key objects in the scene.
Figure 7: The visualization of the selected key points. Red points refer to the key points selected by the proposed method.
Model SemanticKITTI (IoU) % $\uparrow$
0.5m 0.2m 0.1m
LMSCNet [26] 32.23 23.05 3.48
LODE [14] 43.56 47.88 6.06
MID [34] 45.02 41.01 16.98
PVD [51] 21.20 7.96 1.44
LiDiff [24] 42.49 33.12 11.02
LiDiff (Refined) [24] 40.71 38.92 24.75
ScoreLiDAR 38.43 25.75 8.34
ScoreLiDAR (Refined) 37.33 29.57 15.63
Table 11: The IoU evaluation results on the SemanticKITTI dataset.
Model KITTI-360 (IoU) % $\uparrow$
0.5m 0.2m 0.1m
LMSCNet [26] 25.46 16.35 2.99
LODE [14] 42.08 42.63 5.85
MID [34] 44.11 36.38 15.84
LiDiff [24] 42.22 32.25 10.80
LiDiff (Refined) [24] 40.82 36.08 21.34
ScoreLiDAR 36.82 25.49 9.70
ScoreLiDAR (Refined) 33.29 28.60 15.95
Table 12: The IoU evaluation results on the KITTI-360 dataset.
D.6 Experiments on scene occupancy
We calculate the Intersection-Over-Union (IoU) [31] to evaluate the occupancy of the completed scene compared with the ground truth scene. IoU represents the degree of overlap between the voxels in the completed scene and those in the ground truth scene. A higher IoU value indicates a higher completeness of the completed scene. During the evaluation, we considered three different voxel resolutions: $0.5m$, $0.2m$, and $0.1m$. The smaller the voxel resolution, the more fine-grained details are considered in the evaluation metrics, and vice versa.
Tab. 11 and Tab. 12 show the IoU of ScoreLiDAR and existing models. Under low voxel resolutions, ScoreLiDAR achieves comparable IoU values, meaning ScoreLiDAR generates dense and accurate point clouds. When the voxel resolutions become higher, the performance of ScoreLiDAR declines. This is because the existing method is mainly based on signed distance fields, which implement the scene completion using a voxel representation. ScoreLiDAR is point-level scene completion with the input of point clouds obtained from LiDAR scans, which works better at smaller voxel resolutions. The results in Tab. 11 and Tab. 12 do not align with our experimental results and user study findings. Therefore, these results are provided for reference only.
Appendix E Introduction on utilized methods
E.1 Variational score distillation
Variational Score Distillation (VSD), proposed by ProlificDreamer [37], is designed to leverage a pre-trained diffusion model to train a NeRF [22], enabling the rendering of high-quality 3D objects.
Given a text prompt $y$, the probabilistic distribution of all possible 3D representations can be modeled as a probabilistic density $\mu(\theta\|y)$ by a NeRF model parameterized by $\theta$. Let $q_{0}^{\mu}(\boldsymbol{x}_{0}\|c,y)$ as the distribution of the rendered image $\boldsymbol{x}_{0}$ of NeRF given the camera $c$, and $p_{0}(\boldsymbol{x}_{0}\|y)$ as the distribution of the pre-trained text-to-image diffusion model at $t=0$. To generate high-quality 3D objects, ProlificDreamer [37] optimizes the distribution of $\mu$ by minimizing the following KL divergence
$\min_{\mu}D_{\mathrm{KL}}\left(q_{0}^{\mu}\left(\boldsymbol{x}_{0}\mid y\right)\|p_{0}\left(\boldsymbol{x}_{0}\mid y\right)\right)$ (21)
However, directly solving this variational inference problem is challenging because $p_{0}$ is complex, and its high-density regions may be extremely sparse in high-dimensional spaces. Therefore, ProlificDreamer reformulates it as an optimization problem at different time steps $t$, referring to these problems as Variational Score Distillation (VSD),
$\min_{\mu}\mathbb{E}_{t,c}\left[\left(\sigma_{t}/\alpha_{t}\right)\omega(t)D_{\mathrm{KL}}\left(q_{t}^{\mu}\left(\boldsymbol{x}_{t}\mid c,y\right)\|p_{t}\left(\boldsymbol{x}_{t}\mid y\right)\right)\right]$ (22)
Theorem 1 in [37] proves that introducing the additional $t$ does not affect the global optimum of Eq. 21. Theorem 2 in [37] provides the method for optimizing the problem in Eq. 22.
$\frac{\mathrm{d}\theta_{\tau}}{\mathrm{d}\tau}$ $=-\mathbb{E}_{t,\epsilon,c}[\omega(t)(\underbrace{-\sigma_{t}\nabla_{\boldsymbol{x}_{t}}\log p_{t}\left(\boldsymbol{x}_{t}\mid y\right)}_{\text{score of noisy real images }}$ (23)
$-\underbrace{\left(-\sigma_{t}\nabla_{\boldsymbol{x}_{t}}\log q_{t}^{\mu_{\tau}}\left(\boldsymbol{x}_{t}\mid c,y\right)\right)}_{\text{score ofnoisy rendered images }})\frac{\partial\boldsymbol{g}\left(\theta_{\tau},c\right)}{\partial\theta_{\tau}}]$
Here the score of noisy real images is approximated by the pre-trained diffusion model $\boldsymbol{\epsilon}_{pretrain}(\boldsymbol{x}_{t},t,y)$ and the score of noisy rendered images is approximated by another diffusion model $\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},t,c,y)$, which is trained on the rendered images with the standard diffusion objective.
$\min_{\phi}\sum_{i=1}^{n}\mathbb{E}_{t,\epsilon,c}\left[\left\|\boldsymbol{\epsilon}_{\phi}\left(\alpha_{t}\boldsymbol{g}\left(\theta^{(i)},c\right)+\sigma_{t}\boldsymbol{\epsilon},t,c,y\right)-\boldsymbol{\epsilon}\right\|_{2}^{2}\right]$ (24)
In practice, $\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},t,c,y)$ is parameterized by a small UNet or the Low-rank adaptation (LoRA) [10] of the teacher model. With the alternating training of NeRF and $\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},t,c,y)$, ProlificDreamer [37] is ultimately able to generate high-quality 3D objects.
E.2 MinkowskiEngine
Sparse tensor computation plays a critical role in fields such as 3D point cloud processing, computer vision, and physical simulations. Unlike dense tensors, sparse tensors contain a high proportion of zero values and directly applying traditional tensor operations can lead to inefficient use of computational resources. Minkowski Engine [5] addresses these challenges by providing a high-performance framework tailored for sparse tensor computation, enabling efficient operations on high-dimensional sparse data. In this paper, we used the Minkowski Engine to process sparse point cloud data.
Minkowski Engine introduces several innovative approaches to sparse tensor processing.
  • Efficient Sparse Tensor Representation. Sparse tensors are represented using coordinate-value pairs, eliminating the need to store zeros. This representation reduces both memory usage and computational overhead.
  • Sparse Convolution Operations The framework supports high-dimensional sparse convolutions, with kernels designed to adapt to varying sparsity patterns. Optimized memory access patterns and parallel computation strategies ensure high efficiency.
  • Fast Coordinate Mapping Minkowski Engine employs hash tables for rapid coordinate mapping, which accelerates tensor indexing and sparse pattern matching.
  • Automatic Differentiation Support The framework includes built-in support for automatic differentiation, facilitating the training of machine learning models based on sparse tensors.
  • Multi-Dimensional Capability Minkowski Engine can handle sparse tensors of arbitrary dimensions, making it suitable for a wide range of applications, from 2D image processing to 5D simulations.
Minkowski Engine has been widely adopted in various domains including 3D point cloud processing, physical simulations and medical imaging. By significantly improving computational efficiency and scalability, the Minkowski Engine has become a preferred choice for handling sparse tensor computations in both research and industrial applications.
Appendix F Ethical statement
The potential ethical impact of our work is about fairness. As “human” is included as a kind of object in the LiDAR scene, when performing scene completion, it may be necessary to complete human figures. Human-related objects may have data bias related to fairness issues, such as the bias to gender or skin colour. Such bias can be captured by the student model in the training.
F.1 Notification to human subjects
In our user study, we present the notification to subjects to inform the collection and use of data before the experiments.
Dear volunteers, we would like to thank you for supporting our study. We propose ScoreLiDAR, a novel distillation method tailored for 3D LiDAR scene completion, which introduces a structural loss to help the student model capture the geometric structure information. All information about your participation in the study will appear in the study record. All information will be processed and stored according to the local law and policy on privacy. Your name will not appear in the final report. Only an individual number assigned to you is mentioned when referring to the data you provided.
We respect your decision whether you want to be a volunteer for the study. If you decide to participate in the study, you can sign this informed consent form.
The Institutional Review Board approved the use of users’ data of the main authors’ affiliation.
Appendix G Failure examples
Fig. 10 presents some failure cases of ScoreLiDAR. From these examples, it can be observed that ScoreLiDAR exhibits over-completion to some extent, where regions that do not exist are completed. Before the completion, as mentioned in Sec.3 in the main paper, the number of points of the input sparse scan $\mathcal{P}$ is increased by concatenating its points $K$ times and the dense input $\mathcal{P}^{*}$ is obtained. If the number of points of $\mathcal{P}^{*}$ exceeds the actual number of points in the ground truth, it can lead to redundant points in the completed scene. These redundant points may be distributed in areas that do not require completion, resulting in the situations observed in the failure cases.
Figure 8: Completed samples of ScoreLiDAR from KITTI-360 dataset.
Figure 9: Completed samples of ScoreLiDAR from SemanticKITTI dataset.
Figure 10: Failure examples of ScoreLiDAR.